# Embedding in Recommender Systems: A Survey

### v2

@@ 4.3. DHE: Dense Hash Embeddings
In contrast to the aforementioned methods, Google’s Dense Hash Embedding (DHE) (Kang et al., 2020a) introduces several significant innovations. This approach focuses on enhancing hash functions to generate denser embedding vectors. Notably, DHE deploys a DNN-based decoding layer in lieu of conventional embedding tables, leading to markedly superior performance compared to other hash embedding techniques. Remarkably, its performance can even rival that of one-hot encoding. However, it’s worth noting that this enhancement comes at the trade-off of a larger model size relative to other hash embedding methods.
@@
DHE employs a multitude of hash functions, approximately $1,000$ in number, along with a DNN. These hash functions map categorical features like IDs to high-dimensional vectors (e.g., $1,024$ dimensions), with vector elements in the range of $\{1,2,...,m\}$. These vectors can be left normalized, drawn from a uniform distribution, or subjected to transformation into a normal distribution via the Box-Muller method (Box, 1958). Moreover, a DNN with $h$ layers and dropout regularization serves as a decoder. It maps the identity-encoded vectors of dimension $k$ to the final hash embedding vectors of dimension $d$. DHE stands out in its ability to maintain the uniqueness of representations compared to other original hash embedding methods. Furthermore, it addresses the out-of-vocabulary challenge by ensuring the feature embedding is influenced by changes in embedding net parameters. DHE also offers a mechanism for enhancing generalization. This involves concatenating different series of categorical features to the ID-type feature during the decoding process, further improving generalization capabilities.
Finally, We summarize the hash function numbers, hash model size, etc., of all the hash embedding methods mentioned in this section in Table 3.
Table 3. Summary of Hash Embedding Models
Model Embedding Vec ##\#\ #Hash Func ##\#\ #Decoding NN Layer Model Size
Hashing Trick (Weinberger et al., 2009) One-hot 1 1 $O(md)$
Bloom Embed (Serrà and Karatzoglou, 2017) (Multi) One-hot $2\sim 4$ 1 $O(md)$
Hash Embed (Tito Svenstrup et al., 2017) (Multi) One-hot 2 1 $O(nk+md)$
Hybrid Hashing (Zhang et al., 2020) (Multi) One-hot 2 $O(md)$
Q-R Trick (Shi et al., 2020) (Multi) One-hot 2 3 $O(md+\frac{nd^{2}}{m})$
BH (Yan et al., 2021a) (Multi) One-hot & Binary Multi (e.g. 4) $\sim O(\frac{nd}{1000})$
DHE (Kang et al., 2020a) Dense $\sim 1000$ Deep $O(kd_{\rm{NN}}+(h-1)d_{\rm{NN}}^{2}+dd_{\rm{NN}})$
4.4. Survey and Tool
@@
Survey: Kang et al. (Kang et al., 2020a) mention many hashing embedding methods with various techniques in different application scenarios. Ghaemmaghami et al. (Ghaemmaghami et al., 2022) and Li et al. (Li et al., 2023a) present an overview of popular hash embedding models in deep recommendation systems. They mainly discuss the initial purpose of hash functions: compressing the models while controlling the hash collision.
Tool: MurmurHash3101010MurmurHash3: https://github.com/hajimes/mmh3 is a widely used version of the hash function MurMurHash. It can provide $32$-bit and $128$-bit hash. Then CityHash [^11] CityHash: https://github.com/google/cityhash ($64$, $128$, and $256$-bit output) and SpookyHash [^12] SpookyHash: http://burtleburtle.net/bob/hash/spooky.html ($128$-bit output) are both inspired by MurmurHash3 and perform better in terms of compotational efficiency.
4.5. Future Directions
@@
Enhancing Hash Embedding with Multiple Functions In stark contrast to employing a solitary hash function, leveraging multiple hash functions or embracing binary encoding has proven to exert superior control over the scale of the embedding matrix, concurrently reducing the loss of information. Recent innovative designs highlight (Kang et al., 2020a) the beneficial impact of incorporating an augmented number of hash functions. Such a framework possesses the potential to seamlessly replace the extensive and costly hash embedding tables utilized in preceding methodologies, presenting a notably more efficient storage alternative.
@@ 5. AutoML & Embedding
Auto machine learning (AutoML) is a process that automatically generates optimal machine learning solutions for those tasks that are repetitive and time-consuming. In the embedding learning scenario, it could search for the best size of the embedding layer. Concretely, the traditional deep learning approach employs a fixed, uniform embedding size for all features, disregarding the varying importance of each feature in recommendation tasks, which inevitably leads to suboptimal performance. Additionally, the use of excessively large embedding sizes contributes to inflated storage usage and elevated computational expenses. To counter these issues, the AutoML framework introduces the concept of tailoring appropriate embedding sizes for each feature. Neural Architecture Search (NAS) plays a pivotal role in addressing this concern within the realm of AutoML. NAS’s substantial advancements now enable efficient exploration of optimal configurations, including embedding sizes, for recommendation systems within a feasible timeframe.
The methods discussed in this section will be classified by the kinds of the search strategy in NAS which contains reinforcement learning (Kaelbling et al., 1996), gradient-based algorithm (Ruder, 2016), evolutionary algorithm (Qin et al., 2008). In addition, considering that there are a few papers on the evolutionary algorithm, we introduce this category with other unique methods together.
@@ 5.1. Reinforcement learning based method
Reinforcement learning is quite simple to grasp. Think of it like improving your game strategy: if a move or tactic leads to a higher score, reinforcing it enhances your performance. This learning approach involves five main parts: the environment (like a game scenario), the controller or player, actions, rewards, and status. In the case of embedding dimension search (EDS), consider a recommendation model as the setting, evaluating performance to determine rewards. The player, often represented by a policy network, decides what move to make. Rewards are based on how well the model performs. The player uses these rewards to decide how to adjust the current embedding size. The status refers to the player’s situation, like the network parameters.
NIS (Joglekar et al., 2020) initiates Embedding Dimension Search (EDS) by organizing features based on frequency, creating an embedding $\mathbf{E}$ that assigns larger sizes to frequent features. It divides $\mathbf{E}$ into matrices using a fixed search grid. Within this reinforcement learning framework, a recommendation model computes rewards, and a neural network-based player selects actions through soft-max probabilities. The reward integrates memory costs and target objectives into a composite loss function. Asynchronous Advantage Actor-critic (A3C) (Mnih et al., 2016) trains the player and recommendation model alternately. ESAPN (Liu et al., 2020c) dynamically searches for suitable embedding sizes for users and items in stream recommendation. At first, It establishes potential embedding size ranges. Using a recommendation model tailored for stream recommendations, it employs a reinforcement learning approach. The controller, comprising user and item multilayer perceptrons, uses frequency and current embedding size as inputs for the policy network. The controller’s action decides whether to increase or maintain embedding sizes. The reward reflects its predictive capability and guides its decision-making process, which can be calculated from losses of the stream recommendation model as follows:
@@
where the input is a loss sequence of users or items with length $T$, $L^{(u/i)}_{t}$ represents the $t$-th prediction loss of user $u$ or item $i$, and $L$ is the loss of the current item or user. The specific loss function is suitable for the typical recommendation task. To optimize the parameter of the policy network according to the reward, the bell-man equation is formulated as follows:
@@
where $\Phi$ is the parameter of the policy network for the user or item, $a$, and $s$ denotes the action and state, respectively. However, it is hard to calculate its gradient in practice, thus it uses Monte-Carlo sampling to evaluate the $\nabla_{\Phi}J(\Phi)$. In the end, inspired by ENAS, it uses sampled validation data to optimize the policy network and train the recommendation model based on training data.
In summary, reinforcement algorithm-based methods solve the non-differentiable problem of the hard selection in discrete search space and are an efficient and effective search strategy to significantly reduce the search space in embedding size. However, the controller takes a hard selection in a pre-defined embedding candidate size which is likely to achieve a sub-optimal performance. To get higher performance in embedding size selection, the soft selection strategy is widely adopted in gradient-based methods.
@@ 5.2. Gradient-based Method
With the introduction of the Differentiable Neural Architecture Search (DARTS) method proposed by Liu et al. (Liu et al., 2018), gradient-based methods have gained attention. These methods enable the transformation of the search space from discrete to continuous, and the optimization process is guided by gradient calculations or inspired by DARTS. For example, DNIS (Cheng et al., 2020) argues that a pre-defined discrete search space could poor the flexibility of the searching model. Thus, it optimizes the NIS approach by leveraging a soft selection layer to relax the search space to continuous space for performance improvement.
To be specific, it denotes an embedding $E$ with max dimension K by a multi-hot embedding index vectors set $\mathcal{D}=\{d_{1},\cdot\cdot\cdot,d_{N}\}$ and a values vectors set $\mathcal{V}=\{v_{1},\cdot\cdot\cdot,v_{K}\}$ with the search space of $2^{NK}$ size. As same as the NIS, it divides the features into $S$ blocks based on the descending order of frequency. Thus, the search space is reduced from $2^{NK}$ to $2^{SK}$. After that, it defines the soft selection layer as a numerical matrix $\alpha\in\mathbb{R}^{S\times K}$ with elements in the range $[0,1]$, which aims to shrink the elements in embedding within a small number. Subsequently, the approach defines the soft selection layer as a numerical matrix $\mathbf{A}\in\mathbb{R}^{S\times K}$, where the elements are constrained to the range of $[0,1]$. The purpose of this layer is to reduce the values within the embedding to a smaller range. The soft selection operation can be calculated as follow:
@@
where $\widetilde{e}_{i}$ is the $i$-th row of output embedding and $\alpha_{i}$ is $i$-th row of $\mathbf{A}$ matrix. In addition, matrix $\mathbf{A}$ is gained through learning. By inserting a soft selection layer between the embedding layer and interaction layer, it could be optimized jointly like the DARTS method. In particular, it defines a bi-level optimization problem that has been solved by DARTS. Similar to DNIS, autoSrh (Wei et al., 2021) follows the pipeline of DNIS to make tabular Data prediction. The difference is that it debates the idea that setting a pruning threshold could satisfy the requirement of limited memory usage.
@@
Moreover, to improve the hard selection in ESAPN, AutoEMB  (Zhaok et al., 2021) also designed the soft selection layer as a weighted sum operation. Concretely, similar to ESAPN, it defines some candidate embedding sizes for users and items, and employs matrices to transform user and item embedding of candidate sizes into a uniform embedding size for plugging into deep learning-based recommendation models. However, the author noted that thesimple linear transformation process results in significant variation in the values of embeddings with different sizes, rendering them incomparable. To address this issue, the approach applies Batch-Norm with Tanh activation to normalize the transformed user and item embedding vectors. Subsequently, it introduces two multiple perceptron networks to select the embedding size. For an end-to-end differentiable framework, it takes a soft selection layer to calculate the final embedding size.
@@
However, there is another approach called AutoDim (Zhao et al., 2021) that addresses the issue of non-differentiability by employing Gumbel-softmax tricks as a soft selection layer. AutoDim builds upon AutoEMB by extending its input from just users and items to various feature fields such as gender, age, and more. This expansion allows for broader and more comprehensive embeddings that can capture additional contextual information for improved performance.
@@
Another gradient-based idea to find the optimal embedding size is to prune the original embedding through a mask matrix which is learned from train data. For example, AMTL (Yan et al., 2021b) utilizes an Adaptively-Masked Twins-based layer (AMTL) to generate the mask vector. Specifically, the input of AMTL is the frequency attribute of features such that the rank of frequency in a given feature field and AMTL consists of two multiple perceptrons (i.e., h-AML and l-AML) for high-frequency and low-frequency data respectively to avoid the parameter update dominated by high-frequency sample. Then, the final output is the weighted sum to fuse the output of two branch AMLs. Lastly, the approach applies a differentiable temperature softmax function to the output, generating a probability mask vector for selecting an embedding size for high or low-frequency features. However, considering the high training costs of AMTL, SSEDS (Qu et al., 2022) introduces a saliency criterion to determine the significance of each element in the embedding matrix. The saliency score is obtained through a single forward-backward propagation. Using the saliency score, the mask matrix is generated by retaining the top-K important parameters in the embedding, taking into account memory constraints.
@@ 5.3. Other methods
Another unique thought regards the embedding matrix size selection as a regularization problem. Inspired by Soft Threshold Reparameterzation (Kusupati et al., 2020), different from AMTL, PEP (Liu et al., 2021b) directly prune the embedding matrix $\mathbf{E}$ by learning a threshold automatically. Before training the recommendation model, it is argued, according to the Lottery Ticket hypothesis, that the parameters can be initialized by performing an element-wise product with ${\mathbf{E}}_{0}$ and $m$, where $m\in\{0,1\}^{N\times D}$ represents the mask matrix generated from $\hat{\mathbf{E}}$, and ${\mathbf{E}}_{0}$ denotes the unpruned embedding parameters.
@@
There is also a research line of combining learnable anchor embedding matrices to form an optimal embedding matrix. ANT (Liang et al., 2020) and autoDis (Guo et al., 2021a) follow this idea to search for a suitable embedding size. ANT firstly learn meta embedding $A\in\mathbb{R}^{|A|\times d}$ containing a set of anchor embedding $A=\{a_{1},\cdot\cdot\cdot,a_{|A|}\},|A|<<|V|$, where $|V|$ is the number of rows in original embedding $E$. Then, it integrates the set of anchor embedding by using a trainable sparse transformation from $A$ to $E$. However, different from the hard selection of anchor embedding in ANT, autoDis designs a differentiable automatic discretization network to execute a soft selection of meta-embeddings (i.e., anchor embeddings in ANT).
Table 4. Summary of AutoML Embedding, N, K represents # of features and the original length of the feature vector. S and T denote # of the block of the feature and the feature vector, and $\mathbf{F}_{G}$ denotes # of the feature group. U and I indicate # of user and item samples.
Search Strategy Methods Search Space Search space Continuity Storage space Limitation
Reinforcement learning NIS (Joglekar et al., 2020) ${(S+1)}^{T}$ Discrete
ESAPN (Liu et al., 2020c) $2^{(U+I)T}$ Discrete
Gradient-based DNIS (Cheng et al., 2020) $2^{SK}$ Continuous
AutoEMB (Zhaok et al., 2021) $2^{(U+I)T}$ Continuous
AutoDim (Zhao et al., 2021) $T^{|\mathbf{F}_{G}|}$ Continuous
AMTL (Yan et al., 2021b) - Continuous
AutoSrh (Kong et al., 2022) $2^{SK}$ Continuous
SSEDS (Qu et al., 2022) $2^{NK}$ Continuous
Evolutionary algorithm RULE (Chen et al., 2021) $(2^{N}-1)^{|\mathbf{F}_{G}|}$ Discrete
Parameter Regularization PEP (Liu et al., 2021b) $2^{NK}$ Continuous
Anchor embeddings combination ANT (Liang et al., 2020) $2^{SK}$ Continuous
AutoDis (Guo et al., 2021a) $2^{SK}$ Continuous
5.4. Survey and Tool
Survey: In terms of Deep Recommender Systems (DRS), Zheng et al. (Zheng et al., 2022) and Chen et al. (Chen et al., 2022b) both demonstrate how to perform the autoML technique on each component of DRS including feature selection, feature embedding, and feature interaction with a different taxonomy of method.

### v3

@@ 5.3. DHE: Dense Hash Embeddings
In contrast to the aforementioned methods, Google’s Dense Hash Embedding (DHE) (Kang et al., 2020a) introduces several significant innovations. This approach focuses on enhancing hash functions to generate denser embedding vectors. Notably, DHE deploys a DNN-based decoding layer in lieu of conventional embedding tables, leading to markedly superior performance compared to other hash embedding techniques. Remarkably, its performance can even rival that of one-hot encoding. However, this enhancement comes at the trade-off of a larger model size relative to other hash embedding methods.
@@
DHE employs a multitude of hash functions, approximately $1,000$ in number, along with a DNN. These hash functions map categorical features like IDs to high-dimensional vectors (e.g., $1,024$ dimensions), with vector elements in the range of $\{1,2,...,m\}$. These vectors can be left normalized, drawn from a uniform distribution, or subjected to transformation into a normal distribution via the Box-Muller method (Box, 1958). Moreover, a DNN with $h$ layers and dropout regularization serves as a decoder. It maps the identity-encoded vectors of dimension $k$ to final hash embedding vectors of dimension $d$. DHE stands out in its ability to maintain the uniqueness of representations compared to other original hash embedding methods. Furthermore, it addresses the out-of-vocabulary challenge by ensuring the feature embedding is influenced by changes in embedding net parameters. DHE also offers a mechanism for enhancing generalization. This involves concatenating different series of categorical features to the ID-type feature during the decoding process, further improving generalization capabilities.
Building on the dense hash embedding concept, recent work has explored dynamic approaches to hash embedding that can adapt to changing data distributions. Dynamic Sparse Learning (DSL) methods dynamically adjust the sparsity distribution of hash embeddings through pruning and growth strategies, allowing the hash structure to evolve during training (Wang et al., 2024b). This adaptive approach addresses a key limitation of traditional hash embedding methods: their static nature, which cannot respond to shifts in user behavior patterns or item popularity distributions.
5.4. Surveys and Future Directions
@@
Survey. Kang et al. (Kang et al., 2020a) mention many hashing embedding methods with various techniques in different application scenarios. Ghaemmaghami et al. (Ghaemmaghami et al., 2022) present an overview of popular hash embedding models in deep recommendation systems. It primarily discusses the initial purpose of hash functions: compressing models while controlling hash collisions.
@@
Future Direction: Enhancing Hash Embedding with Multiple Functions. In stark contrast to employing a solitary hash function, leveraging multiple hash functions or embracing binary encoding has proven to exert superior control over the scale of the embedding matrix, concurrently reducing the loss of information. Recent innovative designs highlight (Kang et al., 2020a) the beneficial impact of incorporating an augmented number of hash functions. Such a framework possesses the potential to seamlessly replace the extensive and costly hash embedding tables utilized in preceding methodologies, presenting a notably more efficient storage alternative.
@@ 6. AutoML & Embedding
Auto machine learning (AutoML) is a process that automatically generates optimal machine learning solutions for repetitive and time-consuming tasks. In the embedding learning scenario, it could search for the best size of the embedding layer. Concretely, the traditional deep learning approach employs a fixed, uniform embedding size for all features, disregarding the varying importance of each feature in recommendation tasks, which inevitably leads to suboptimal performance. Additionally, the use of excessively large embedding sizes contributes to inflated storage usage and elevated computational expenses. To counter these issues, the AutoML framework introduces the concept of tailoring appropriate embedding sizes for each feature. Neural Architecture Search (NAS) plays a pivotal role in addressing this concern within the realm of AutoML. NAS’s substantial advancements now enable efficient exploration of optimal configurations, including embedding sizes, for recommendation systems within a feasible timeframe.
The methods discussed in this section will be classified by the kinds of search strategies in NAS, which include reinforcement learning (Kaelbling et al., 1996), gradient-based algorithm (Ruder, 2016), and evolutionary algorithm (Qin et al., 2008). In addition, considering that there are a few papers on the evolutionary algorithm, we introduce this category with other unique methods together. These search strategies and their corresponding models are detailed in Table 6.
Table 6.
AutoML & Embedding
AutoML & Embedding Reinforcement learning based method (§6.1) Key Models: NIS (Joglekar et al., 2020), ESAPN (Liu et al., 2020c)
Gradient-based Method (§6.2) Soft Selection Models: DNIS (Cheng et al., 2020), autoSrh (Wei et al., 2021), AutoEMB (Zhaok et al., 2021), AutoDim (Zhao et al., 2021b)
Pruning-based Models: AMTL (Yan et al., 2021b), SSEDS (Qu et al., 2022)
Budget-constrained Models: BET (Qu et al., 2024)
Other methods (§6.3) Evolutionary Algorithms: RULE (Chen et al., 2021)
Regularization/Pruning: PEP (Liu et al., 2021b)
Anchor Embeddings: ANT (Liang et al., 2020), autoDis (Guo et al., 2021a)
@@ 6.1. Reinforcement learning based method
Reinforcement learning is quite simple to grasp. Think of it like improving your game strategy: if a move or tactic leads to a higher score, reinforcing it enhances your performance. This learning approach involves five main parts: the environment (like a game scenario), the controller or player, actions, rewards, and status. In the case of embedding dimension search (EDS), consider a recommendation model as the setting, evaluating performance to determine rewards. The player, often represented by a policy network, decides what move to make. Rewards are based on how well the model performs. The player uses these rewards to decide how to adjust the current embedding size. The status refers to the player’s situation, like the network parameters. The entire process, often referred to as Embedding Dimension Search (EDS), is illustrated in Fig. 6.
Figure 6. The reinforcement learning loop for Embedding Dimension Search (EDS). The Controller (Agent), based on the current state St, selects an action at (e.g., modifying embedding dimensions). The Recommendation Model (Environment) executes this action and returns a reward Rt and the next state St+1. The agent learns by optimizing its policy to maximize future rewards.
NIS (Joglekar et al., 2020) initiates Embedding Dimension Search (EDS) by organizing features based on frequency, creating an embedding $\mathbf{E}$ that assigns larger sizes to frequent features. It divides $\mathbf{E}$ into matrices using a fixed search grid. Within this reinforcement learning framework, a recommendation model computes rewards, and a neural network-based player selects actions through softmax probabilities. The reward integrates memory costs and target objectives into a composite loss function. Asynchronous Advantage Actor-critic (A3C) (Mnih et al., 2016) trains the player and recommendation model alternately. ESAPN (Liu et al., 2020c) dynamically searches for suitable embedding sizes for users and items in stream recommendation. At first, It establishes potential embedding size ranges. Using a recommendation model tailored for stream recommendations, it employs a reinforcement learning approach. The controller, comprising user and item multilayer perceptrons, uses frequency and current embedding size as inputs for the policy network. The controller’s action decides whether to increase or maintain embedding sizes. The reward reflects its predictive capability and guides its decision-making process, which can be calculated from losses of the stream recommendation model as follows:
@@
where the input is a loss sequence of users or items with length $T$, $L^{(u/i)}_{t}$ represents the $t$-th prediction loss of user $u$ or item $i$, and $L$ is the loss of the current item or user. The specific loss function is suitable for the typical recommendation task. To optimize the parameters of the policy network according to the reward, the Bellman equation is formulated as follows:
@@
where $\Phi$ is the parameter of the policy network for the user or item, $a$, and $s$ denote the action and state, respectively. However, it is hard to calculate its gradient in practice; thus, it uses Monte-Carlo sampling to evaluate the $\nabla_{\Phi}J(\Phi)$. In the end, inspired by ENAS, it uses sampled validation data to optimize the policy network and train the recommendation model based on training data.
In summary, reinforcement algorithm-based methods solve the non-differentiable problem of the hard selection in a discrete search space and are an efficient and effective search strategy to significantly reduce the search space in embedding size. However, the controller takes a hard selection in a pre-defined embedding candidate size, which is likely to achieve a sub-optimal performance. To get higher performance in embedding size selection, the soft selection strategy is widely adopted in gradient-based methods.
@@ 6.2. Gradient-based Method
Gradient-based methods, such as the Differentiable Neural Architecture Search (DARTS) method proposed by Liu et al. (Liu et al., 2018), have gained attention. These methods transform the search space from discrete to continuous, and the optimization process is guided by gradient calculations or inspired by DARTS. For example, DNIS (Cheng et al., 2020) argues that a pre-defined discrete search space could limit the flexibility of the searching model. Thus, it optimizes the NIS approach by leveraging a soft selection layer to relax the search space to continuous space for performance improvement.
Specifically, it denotes an embedding $E$ with max dimension K by a multi-hot embedding index vectors set $\mathcal{D}=\{d_{1},\cdot\cdot\cdot,d_{N}\}$ and a values vectors set $\mathcal{V}=\{v_{1},\cdot\cdot\cdot,v_{K}\}$ with the search space of $2^{NK}$. As same as the NIS, it divides the features into $S$ blocks based on the descending order of frequency. Thus, the search space is reduced from $2^{NK}$ to $2^{SK}$. After that, it defines the soft selection layer as a numerical matrix $\alpha\in\mathbb{R}^{S\times K}$ with elements in the range $[0,1]$, which aims to shrink the elements in embedding within a small number. The purpose of this layer is to reduce the values within the embedding to a smaller range. The soft selection operation can be calculated as follows:
@@
where $\widetilde{e}_{i}$ is the $i$-th row of output embedding and $\alpha_{i}$ is $i$-th row of $\mathbf{A}$ matrix. In addition, matrix $\mathbf{A}$ is gained through learning. By inserting a soft selection layer between the embedding layer and interaction layer, it could be optimized jointly like the DARTS method. In particular, it defines a bi-level optimization problem that has been solved by DARTS.
Similar to DNIS, autoSrh (Wei et al., 2021) follows the pipeline of DNIS to make tabular Data prediction. The difference is that it debates the idea that setting a pruning threshold could satisfy the requirement of limited memory usage.
@@
Moreover, to improve the hard selection in ESAPN, AutoEMB  (Zhaok et al., 2021) proposed the soft selection layer as a weighted sum operation. Concretely, similar to ESAPN, it defines some candidate embedding sizes for users and items, and employs matrices to transform user and item embedding of candidate sizes into a uniform embedding size for plugging into deep learning-based recommendation models. However, the author noted that thesimple linear transformation process results in significant variation in the values of embeddings with different sizes, rendering them incomparable. To address this issue, the approach applies Batch-Norm with Tanh activation to normalize the transformed user and item embedding vectors. Subsequently, it introduces two multiple perceptron networks to select the embedding size. For an end-to-end differentiable framework, it takes a selection layer to get the embedding size.
@@
However, there is another approach called AutoDim (Zhao et al., 2021b) that addresses the issue of non-differentiability by employing Gumbel-softmax tricks as a soft selection layer. AutoDim builds upon AutoEMB by extending its input from just users and items to various feature fields such as gender, age, and more. This expansion allows for more comprehensive embeddings that can capture additional contextual information for improved performance. The evolution of gradient-based embedding search has led to more sophisticated approaches that address real-world deployment constraints. Budgeted Embedding Table (BET) (Qu et al., 2024), which advances beyond traditional AutoML approaches by incorporating explicit memory budget constraints into the search process. Unlike previous methods that optimize embedding dimensions independently, BET formulates the problem as finding table-level actions that are guaranteed to meet pre-specified memory budgets while maximizing recommendation performance.
@@
Another gradient-based idea to find the optimal embedding size is to prune the original embedding through a mask matrix, which is learned from the training data. For example, AMTL (Yan et al., 2021b) utilizes an Adaptively-Masked Twins-based layer (AMTL) to generate the mask vector. Specifically, the input of AMTL is the frequency attribute of features, such that the rank of frequency in a given feature field and AMTL consists of two multiple perceptrons (i.e., h-AML and l-AML) for high-frequency and low-frequency data, respectively, to avoid the parameter update dominated by high-frequency samples. Then, the final output is the weighted sum to fuse the output of two branch AMLs. Lastly, the approach applies a differentiable temperature softmax function to the output, generating a probability mask vector for selecting an embedding size for high or low-frequency features. However, considering the high training costs of AMTL, SSEDS (Qu et al., 2022) introduces a saliency criterion to determine the significance of each element in the embedding matrix. The saliency score is obtained through a single forward-backward propagation. Using the saliency score, the mask matrix is generated by retaining the top-K important parameters, taking into account memory constraints.
@@ 6.3. Other methods
Another novel perspective views embedding matrix size selection as a regularization problem. Inspired by Soft Threshold Reparameterization (Kusupati et al., 2020), different from AMTL, PEP (Liu et al., 2021b) directly prunes the embedding matrix $\mathbf{E}$ by learning a threshold. Before training the recommendation model, it is argued, according to the Lottery Ticket hypothesis, that the parameters can be initialized by performing an element-wise product between ${\mathbf{E}}_{0}$ and $m$, where ${\mathbf{E}}_{0}$ denotes the unpruned embedding parameters, and $m\in\{0,1\}^{N\times D}$ represents the mask matrix generated from $\hat{\mathbf{E}}$.
@@
Another line of research combines learnable anchor embedding matrices to form an optimal embedding matrix. ANT (Liang et al., 2020) and autoDis (Guo et al., 2021a) follow this idea to search for a suitable embedding size. ANT first learns a meta embedding $A\in\mathbb{R}^{|A|\times d}$ containing a set of anchor embeddings $A=\{a_{1},\cdot\cdot\cdot,a_{|A|}\},|A|<<|V|$, where $|V|$ is the number of rows in original embedding $E$. Then, it integrates the set of anchor embeddings with a trainable sparse transformation from $A$ to $E$. However, different from the hard selection of anchor embeddings in ANT, autoDis designs a differentiable automatic discretization network to execute a soft selection of meta-embeddings (i.e., anchor embeddings in ANT).
6.4. Surveys and Future Directions